# DUNIA: Pixel-Sized Embeddings via Cross-Modal Alignment for Earth Observation Applications

### v1

@@ 3.3 Multitemporal Image Processing Model
The multitemporal image model (Figure 2) is designed to process temporal sequences of median composite images, following an autoencoder (AE) architecture. The input to this model is a sequence of images $\mathcal{M}\in\mathbb{R}^{T\times H\times W\times C}$ covering the same area as the input to the pre-training model. These images come from $T$ Sentinel- 1 & 2 acquisition dates that overlap with the dates used to generate the composite image $\mathcal{I}$. This design choice, unlike multi-temporal pre-training models that require time series data during training and inference, or mono-temporal models limited to a single date, enables the model to capture some phenological traits without needing time series data during inference.
@@ 3.4 Pre-Training Objective
DUNIA is pre-trained on pixel-waveform and pixel-pixel alignment, alongside modality reconstruction for the modality-specific AEs. For the alignment task, we rely on two, similarly performing, non-negative-pair contrastive losses, namely VICReg (Bardes et al., 2021) and Zero-CL (Zhang et al., 2021). The selection between either loss functions is determined by the number of available elements within a mini-batch and the requirements of each loss function. VICReg requires large batch sizes to perform well (Bardes et al., 2021), which is a non-issue for the pixel-pixel contrastive objective given the high number of pixels available in each mini-batch of images. However, this is not suitable for the waveforms which are few in number within the mini-batch. On the other hand, Zero-CL performs well even for small batch sizes but faces computational bottlenecks with very large batches. Consequently, we adopt Zero-CL for pixel-waveform alignment and VICReg for pixel-pixel alignment. The formulation of these two objectives are detailed below.
@@ 3.4.1 Pixel-Waveform Alignment
Zero-CL replaces the alignment and uniformity terms in negative-pair-based contrastive losses (Arora et al., 2019) with, respectively, an instance-wise contrastive loss ($\mathcal{L}_{\text{Ins}}$) and a feature-wise contrastive loss ($\mathcal{L}_{\text{Fea}}$). The overall loss $\mathcal{L}_{Zero-CL}$ is $\mathcal{L}_{\text{Fea}}+\mathcal{L}_{\text{Ins}}$. $\mathcal{L}_{Zero-CL}$ is applied on $Z^{A}\in\mathbb{R}^{G\times D_{p}}$, the $L_{2}$ normalized pixel embeddings from $\mathcal{O}^{\mathcal{V}}$, and $Z^{B}\in\mathbb{R}^{G\times D_{p}}$ their corresponding $L_{2}$ normalized waveform embeddings from $\mathcal{O}^{\mathcal{W}}$, where $G$ is the number of available GEDI samples in a given mini-batch. The mathematical formulation for this loss can be found in Appendix C.
@@ 3.4.2 Pixel-Pixel Alignment
Following Bardes (2021) we set $\alpha_{v}$ and $\beta_{i}$ to 25.0 and $\gamma_{c}$ to 1.0. The mathematical formulation of VICReg can be found in Appendix C.
@@ 3.5 Waveform Generation
For a detailed description of the diffusion process, including the denoising objective, noise schedule, and latent variable sampling, we refer the reader to Appendix A.4.
@@ 4 Experimental Evaluation
We provide a detailed experimental evaluation showing that DUNIA is able to achieve high performance across a variety of tasks in zero-shot and fine-tuned settings, including land cover classification, crop mapping, and vertical forest structure analysis (cover, height and GEDI waveform retrieval). We leverage diverse datasets to inform the model on horizontal structures–Sentinel- 1 & 2, and vertical structures with GEDI waveforms.
The following section contains a summary of the experimental setup, with full details in the appendix. Model configuration and optimisation can be found in Appendix D. For clarity, during the pre-training phase, we set the input image size to $64\times 64$ pixels, with 14 channels from stacked Sentinel-1 & 2 image composites. The embedding dimension is set to 64 (i.e., $D_{p}=64$). During inference, the input image size can be of any size, but we inferred on inputs of shape $256\times 256$ pixels.
@@ 4.1.1 Datasets
We used Sentinel-2 Level-2A surface reflectance data from Google Earth Engine, including 10 m and 20 m spatial resolution bands with the latter upscaled to 10 m. Two sets of mosaics were created over the entire metropolitan French territory: a single leaf-on season mosaic (April-September 2020) for the pre-trained model, and three four-month mosaics (October 2019-September 2020) for the multitemporal AE. Cloud filtering was applied using the S2 Cloud Probability dataset.
Sentinel-1 data were obtained from the S1A and S1B satellites operating at C-band, collected in Interferometric Wide swath mode with VV and VH polarizations. The data was calibrated, geometrically corrected, and resampled to 10 m resolution. Similar to S2, we created two sets of mosaics with normalized backscattering coefficients.
The Global Ecosystem Dynamics Investigation (GEDI) is a full-waveform LiDAR sensor on the International Space Station (ISS), operational between 51.6 N and 51.6 S from 2019-2023. We used Level 1B, 2A, and 2B data from April 2019 to December 2021, extracting waveforms, geolocation, height metrics ($\mathcal{W}_{rh}$), canopy cover ($\mathcal{W}_{c}$), and Plant Area Index ($\mathcal{W}_{pai}$). After quality filtering following Fayad et al. (2024) and seasonal selection, the dataset contained $\approx$ 19 million waveforms; covering less than 1% of the total surface area of France.
Overall, our pre-training dataset consisted of 836K $64\times 64$ pixels Sentinel- 1 & 2 images with, on average, 26 corresponding GEDI waveforms per image.
@@
We evaluated our model on seven downstream tasks using labels from various data products at resolutions matching or lower than our model’s output (10 m). PureForest ($PF$) provides a benchmark dataset of ground truth patches for classifying mono-specific forests in France, featuring high-resolution imagery and annotations for over 135K $50\times 50$ m (i.e., 5x5 pixels) patches across 13 tree species (Gaydon & Roche, 2024). CLC+Backbone ($CLS_{+}$) is a pan-European land cover inventory for 2021, utilizing Sentinel-2 time series (2020-2022) and a TempCNN classifier (Pelletier et al., 2019) to produce a 10 m raster indicating the dominant land cover among 11 classes. PASTIS is a crop mapping dataset by Garnot et al. (2021), covering 18 crop classes and 1 background class with 2433 densely annotated $128\times 128$ pixels images at 10 m resolution. The Vertical Structure dataset assesses model performance in mapping forest heights, canopy cover, plant area index, and waveform retrieval at 10 m resolution, relying on GEDI-derived products as presented earlier. When available, we used the train/val/test split used by the references (i.e., $PF$ and PASTIS). For the $CLS_{+}$ dataset, we used the same split as the unsupervised dataset, which followed a 65/10/25 split.
@@ 4.1.2 Performance Evaluation
For the low-shot fine-tuning, we froze the entire pre-trained network except for the last two NA layers in each decoder, which were appended with an output head consisting of two sequential $1\times 1$ convolutional layers. The first layer halves the input channels, and the second projects the reduced representation to the desired output size.
@@
We compared DUNIA in the fine-tuned setting to two current state-of-the-art Earth Observation FMs: CROMA (Fuller et al., 2023) and AnySat (Astruc et al., 2024). CROMA takes as input Sentinel- 1 & 2 imagery, while Anysat is pre-trained using Sentinel- 1 & 2 times series as well as very high-resolution imagery. For a fair comparison, we refined the pre-training of both models for 200K steps using our datasets and the training details from the respective papers. For AnySat, we used multi-temporal Sentinel-1 & 2 mosaics with three time steps and included SPOT images at 1.5 m resolution as an additional input modality during pre-training but fine-tuned using only Sentinel- 1 & 2. Both models were evaluated on all downstream tasks except waveform generation, as they do not support this task.
@@ 4.2 Results
Our evaluation shows that embeddings from our proposed framework, combined with simple zero-shot classifiers, often surpass specialized supervised models. Even with minimal labeled data, our model demonstrates strong low-shot performance, rivaling or exceeding state-of-the-art methods.
@@ 4.2.1 Zero-Shot Classification Performance
Table 1: Top-1 retrieval-based zero-shot classification performance of DUNIA for different retrieval settings. SOTA results, when available, represent the best-performing supervised model for any given dataset. M1 are from Schwartz et al. (2023), M1 are retrained models from M1 for that particular dataset. M2 comes from Garnot & Landrieu (2021). M3 comes from the model trained with very high-resolution aerial imagery developed in Gaydon & Roche (2024). and are DUNIA query embeddings from $\mathcal{O^{V}}$ and $\mathcal{O^{H}}$ respectively. S represents the number of samples in the retrieval database, with im meaning a $64\times 64$ pixels fully annotated image, and l meaning a single annotated pixel. KNN represents the K nearest neighbors used in the distance-weighted voting. $\text{KNN}_{b}$ represents the best KNN value for a given dataset. $\mathcal{W}^{**}$ represents performance results for vertical structures higher than 5 m. Best results for each metric are highlighted in bold.
Dataset Metric SOTA S DUNIA
KNN = 50 KNN = 5 KNN =1 10%S, $\text{KNN}_{b}$ 5%S, $\text{KNN}_{b}$
$\mathcal{W}_{rh}$ $rmse$ $(r)$ [M1]    5.2 (.77) 50K l 2.0 (.93) 2.1 (.93) 2.1 (.91) 2.1 (.92) 2.1 (.92)
$\mathcal{W}_{c}$ $rmse$ $(r)$ [M1]    22.1 (.54) 50K l 11.7 (.89) 12.1 (.84) 12.3 (.86) 12.4 (.86) 12.0 (.84)
$\mathcal{W}_{pai}$ $rmse$ $(r)$ [M1]    1.5 (.35) 50K l 0.71 (.75) 0.74 (.73) 0.74 (.72) 0.72 (.75) 0.72 (.74)
$CLC_{+}$ $wF1$ 500 im 80.1% 74.3% 69.5% 75.2% 70.2%
$PASTIS$ $OA$ [M2]    84.2% 500 im 56.2% 54.1% 49.5% 52.4% 48.3%
$PF$ $wF1$ [M3]    74.6% 50K l 73.8% 76.0% 75.8% 73.5% 70.9%
$\mathcal{W}^{**}$ $r$ 50K l .70 .67 .66
This number of samples is considered as a low data regime due to: 1) a total coverage area of $\approx 31Km^{2}$ 2) only $\approx 0.25\%$ of data required compared to supervised models.
Table 2: Finetuning performance of DUNIA and the two competing models (i.e., AnySat (Astruc et al., 2024), and CROMA (Fuller et al., 2023)). and are DUNIA’s embeddings from $\mathcal{O^{V}}$ and $\mathcal{O^{H}}$ respectively. S represents the number of samples used for the finetuning, with im meaning a $64\times 64$ pixels fully annotated image, and l meaning a single annotated pixel. $\mathcal{W}^{**}$ represents performance results for vertical structures higher than 5 m. Best results for each metric are highlighted in bold for the first set of samples configuration and underlined for the second set of samples.
Dataset Metric Samples (S) DUNIA AnySat Croma S DUNIA AnySat Croma
$\mathcal{W}_{rh}$ $rmse$ $(r)$ 5K im 1.34 (.95) 2.75 (.89) 3.52 (.78) 1K im 1.44 (.93) 2.80 (.89) 3.62 (.76)
$\mathcal{W}_{c}$ $rmse$ $(r)$ 5K im 9.8 (.85) 12.1 (.79) 14.2 (.73) 1K im 10.0 (.83) 12.4 (.79) 14.5 (.72)
$\mathcal{W}_{pai}$ $rmse$ $(r)$ 5K im 0.62 (.71) 0.95 (.67) 1.2 (.61) 1K im 0.61 (.70) 0.94 (.67) 1.5 (.60)
$CLC_{+}$ $wF1$ 5K im 90.3% 90.1% 86.4% 1K im 89.4% 89.5% 85.9%
$PASTIS$ $wF1$ 1.5K im 77.0% 81.1% 73.3% 300 im 75.3% 80.2% 71.0%
$PF$ $wF1$ 50K l 82.2% 82.3% 80.5% 10K l 80.1% 80.0% 80.2%
$\mathcal{W}^{**}$ $r$approx19M l .78 $\approx$3.8M l .75
@@
Zero-shot results in Table 1 demonstrate that for vertical structure products ($\mathcal{W}_{rh}$, $\mathcal{W}_{c}$, $\mathcal{W}_{pai}$) and tree species identification ($PF$), DUNIA with S samples and KNN=50 outperforms specialist models. It achieves RMSE (r) improvements of 3.2 m (.16), 10.4% (.35), and 0.79 (.4) for canopy height, cover, and plant area index, and matching the mapping quality of supervised methods (Figure 3). For tree species identification, DUNIA slightly underperforms by 0.8% at (S=20, KNN=50) but exceeds the baseline by 1.4% when decreasing the number of neighbors (KNN=5) used for the distance-weighted voting. For land cover mapping ($CLC_{+}$), DUNIA achieves strong performance with a $wF1$ score of 80.1% (S=500 im, KNN=50). However, for crop type mapping ($PASTIS$), DUNIA significantly underperforms compared to the state-of-the-art, which is likely due to the variability of the phenological cycles of crops that cannot be well captured by a single median composite.
@@
For lower KNN values, the performance of vertical structure-related products and tree species identification remains stable. In contrast, land cover and crop type mapping experience significant performance drops, indicating that the model needs more samples for accurate classification. For retrieval databases with fewer samples, vertical structure-related products ($\mathcal{W}_{rh}$, $\mathcal{W}_{c}$, $\mathcal{W}_{pai}$) maintain their performance and map qualities (Figure 4), while other products ($CLC_{+}$, $PASTIS$, $PF$) show a substantial degradation in performance.
@@
Regarding waveform ($\mathcal{W}$) retrieval performance, we observe that our model is able to retrieve relevant waveforms given a pixel embedding as a query, with a correlation coefficient of .70% for S=50k l, and decreases by a small factor with smaller retrieval database sizes. This is to be expected given the small subset of waveforms in the retrieval database that might not reflect all possible variations for a particular height class. Retrieved vs. reference waveforms can be found in Figures 5, 6, 7 and 8.
4.2.2 Finetuning Performance

### v2

@@ 3.3 Multitemporal Image Processing Model
The multitemporal image model (Figure 2) is designed to process temporal sequences of median composite images, following an autoencoder (AE) architecture. The input to this model is a sequence of images $\mathcal{M}\in\mathbb{R}^{T\times H\times W\times C}$ covering the same area as the input to the pre-training model. These images come from $T$ S-1 & 2 acquisition dates that overlap with the dates used to generate the composite image $\mathcal{I}$. This design choice, unlike multi-temporal pre-training models that require time series data during training and inference, or mono-temporal models limited to a single date, is a trade-off that enables the model to capture some phenological traits without needing time series data during inference. Indeed, while image composites such as a median composite may be richer than single-date images, they are still less informative than a full time series, as they only provide a median reflectance value over a given period. Conversely, even this simple form of aggregation preserves significantly more information than a single-date image.
@@ 3.4 Pre-Training Objective
DUNIA is pre-trained on pixel-waveform and pixel-pixel alignment, alongside modality reconstruction for the modality-specific AEs. For the alignment task, we rely on two, similarly performing, non-negative-pair contrastive losses, namely VICReg (Bardes et al., 2021) and Zero-CL (Zhang et al., 2021). The selection between either loss functions is determined by the number of available elements within a mini-batch and the requirements of each loss function. VICReg requires large batch sizes to perform well (Bardes et al., 2021), which is a non-issue for the pixel-pixel contrastive objective given the high number of pixels available in each mini-batch of images. However, this is not suitable for the waveforms which are few in number within the mini-batch. On the other hand, Zero-CL performs well even for small batch sizes but faces computational bottlenecks with very large batches.
@@ 3.4.1 Pixel-Waveform Alignment
Zero-CL replaces the alignment and uniformity terms in negative-pair-based contrastive losses (Arora et al., 2019) with, respectively, an instance-wise contrastive loss ($\mathcal{L}_{\text{Ins}}$) and a feature-wise contrastive loss ($\mathcal{L}_{\text{Fea}}$). The overall loss $\mathcal{L}_{Zero-CL}$ is $\mathcal{L}_{\text{Fea}}+\mathcal{L}_{\text{Ins}}$. $\mathcal{L}_{Zero-CL}$ is applied on $Z^{A}\in\mathbb{R}^{G\times D_{p}}$, the $L_{2}$ normalized pixel embeddings from $\mathcal{O}^{\mathcal{V}}$, and $Z^{B}\in\mathbb{R}^{G\times D_{p}}$ their corresponding $L_{2}$ normalized waveform embeddings from $\mathcal{O}^{\mathcal{W}}$, where $G$ is the number of available GEDI samples in a given mini-batch. The formulation for this loss can be found in Appendix B.
@@ 3.4.2 Pixel-Pixel Alignment
Following Bardes (2021) we set $\alpha_{v}$ and $\beta_{i}$ to 25.0 and $\gamma_{c}$ to 1.0. VICReg formulation can be found in Appendix B.
@@ 3.5 Waveform Generation
For a detailed description of the diffusion process, including the denoising objective, noise schedule, and latent variable sampling, we refer the reader to Appendix C.
@@ 4 Experimental Evaluation
We provide a detailed experimental evaluation showing that DUNIA is able to achieve high performance across a variety of tasks in zero-shot and fine-tuned settings, including land cover classification, crop mapping, and vertical forest structure analysis (e.g., canopy cover, height and GEDI waveform retrieval). We leverage diverse datasets to inform the model on horizontal structures – S-1 & 2, and vertical structures with GEDI waveforms.
The following section contains a summary of the experimental setup, with full details in the appendix. Model configuration and optimisation can be found in Appendix E. For clarity, during the pre-training phase, we set the input image size to $64\times 64$ pixels, with 14 channels from stacked S-1 & 2 image composites. The embedding dimension is set to 64 (i.e., $D_{p}=64$). During inference, the input image can be of any size, but we inferred on $256\times 256$ pixel images. Our code is available at github.com/AI4Forest/DUNIA.
@@ 4.1.1 Datasets
We used S-2 Level-2A surface reflectance data from Google Earth Engine, including 10 m and 20 m spatial resolution bands with the latter upscaled to 10 m. Two sets of mosaics were created over the entire metropolitan French territory: a single leaf-on season mosaic (April-September 2020) for the pre-trained model, and three four-month mosaics (October 2019-September 2020) for the multitemporal AE. Cloud filtering was applied using the S-2 Cloud Probability dataset.
S-1 data were obtained from the S-1A and S-1B satellites operating at C-band, collected in Interferometric Wide swath mode with VV and VH polarizations. The data was calibrated, geometrically corrected, and resampled to 10 m resolution. Similar to S-2, we created two sets of mosaics with normalized backscattering coefficients.
GEDI is a full-waveform LiDAR sensor on the International Space Station (ISS), operational between 51.6 N and 51.6 S from 2019-2023. We used Level 1B, 2A, and 2B data from April 2019 to December 2021, extracting waveforms, geolocation, height metrics ($\mathcal{W}_{rh}$), canopy cover ($\mathcal{W}_{c}$), and Plant Area Index ($\mathcal{W}_{pai}$). After quality filtering following Fayad et al. (2024) and seasonal selection, the dataset contained $\approx$ 19 million waveforms, covering less than 1% of the total surface area of France.
Overall, our pre-training dataset consisted of 836K $64\times 64$ pixels S-1 & 2 images with, on average, 26 GEDI waveforms per image. Additional details on the pre-training datasets can be found in Appendix F.
@@
We evaluated our model on seven downstream tasks using labels from various data products at resolutions matching or lower than our model’s output (10 m). PureForest ($PF$) provides a benchmark dataset of ground truth patches for classifying mono-specific forests in France, featuring high-resolution imagery and annotations for over 135K $50\times 50$ m (i.e., 5x5 pixels) patches across 13 tree species (Gaydon & Roche, 2024). CLC+Backbone ($CLS_{+}$) is a pan-European land cover inventory for 2021, utilizing S-2 time series (2020-2022) and a TempCNN classifier (Pelletier et al., 2019) to produce a 10 m raster indicating the dominant land cover among 11 classes. PASTIS is a crop mapping dataset by Garnot et al. (2021), covering 18 crop classes and 1 background class with 2433 densely annotated $128\times 128$ pixels images at 10 m resolution. The vertical structure dataset assesses model performance in mapping forest heights, canopy cover, plant area index, and waveform retrieval at 10 m resolution, relying on GEDI-derived products as presented earlier. When available, we used the train/val/test split used by the references (i.e., $PF$ and PASTIS). For the $CLS_{+}$ dataset, we used the same split as the unsupervised dataset, which followed a 65/10/25 split.
@@ 4.1.2 Performance Evaluation
For the low-shot fine-tuning, we froze the entire pre-trained network except for the last NA layer in each decoder, which were appended with an output head consisting of two sequential $1\times 1$ convolutional layers. The first layer halves the input channels, and the second projects the reduced representation to the desired output size.
@@
We compared DUNIA in the fine-tuned setting to five current state-of-the-art Earth Observation FMs: SatMAE (Cong et al., 2022), DOFA (Xiong et al., 2024), DeCUR (Wang et al., 2024), CROMA (Fuller et al., 2023) and AnySat (Astruc et al., 2024). SatMAE takes as input S-2 imagery, DOFA, DeCUR, and CROMA take as input S-1 & 2 imagery, while Anysat is pre-trained using S-1 & 2 times series as well as very high-resolution imagery. For a fair comparison, we pre-trained all competing models (using pre-trained weights when available) for 250K steps using our datasets and the training details from the respective papers. For AnySat, we used multi-temporal S-1 & 2 mosaics with three time steps and included SPOT images at 1.5 m resolution as an additional input modality during pre-training but fine-tuned using only S-1 & 2. All competing models were evaluated on all downstream tasks except waveform generation, as they do not support this task.
Table 1: Top-1 retrieval-based zero-shot classification performance of DUNIA for different retrieval settings. SOTA results, when available, represent the best-performing supervised models for any given dataset. are retrained models for a particular dataset. and are DUNIA query embeddings from $\mathcal{O^{V}}$ and $\mathcal{O^{H}}$ respectively. S represents the number of samples in the retrieval database, with im meaning a $64\times 64$ pixels fully annotated image, and l meaning a single annotated pixel. KNN represents the $k$ nearest neighbors used in the distance-weighted voting. $\mathcal{W}^{**}$ represents performance results for vertical structures higher than 5 m. Best scores are in bold.
Dataset Metric SOTA S DUNIA
KNN = 50 KNN = 5 KNN =1
Schwartz et al. (2023)    5.2 (.77)
Liu et al. (2023)    5.2 (.76)
Tolan et al. (2024)    8.5 (.59)
$\mathcal{W}_{rh}$ RMSE (r) Lang et al. (2023)    5.6 (.76) 50K l 2.0 (.93) 2.1 (.93) 2.1 (.91)
$\mathcal{W}_{c}$ RMSE (r) Schwartz et al. (2023)    22.1 (.54) 50K l 11.7 (.89) 12.1 (.84) 12.3 (.86)
$\mathcal{W}_{pai}$ RMSE (r) Schwartz et al. (2023)    1.5 (.35) 50K l 0.71 (.75) 0.74 (.73) 0.74 (.72)
$CLC_{+}$ wF1 500 im 80.1 74.3 69.5
$PASTIS$ OA Garnot & Landrieu (2021)    84.2 500 im 56.2 54.1 49.5
$PF$ wF1 Gaydon & Roche (2024)    74.6 50K l 73.8 76.0 75.8
$\mathcal{W}^{**}$ r 50K l .70
This number of samples is considered as a low data regime due to: 1) a total coverage area of $\approx 31Km^{2}$ 2) only $\approx 0.25\%$ of data required compared to some supervised models.
@@ 4.2 Results
Our evaluation shows that embeddings from our proposed framework, combined with simple zero-shot classifiers, surpass specialized supervised models in several tasks. In the fine-tuned setting, our model demonstrates strong low-shot performance, rivaling or exceeding state-of-the-art methods.
@@ 4.2.1 Zero-Shot Classification Performance
Zero-shot results in Table 1 demonstrate that for vertical structure products ($\mathcal{W}_{rh}$, $\mathcal{W}_{c}$, $\mathcal{W}_{pai}$), DUNIA with KNN=50 outperforms specialist models. It achieves RMSE (r) improvements of 3.2 m (.16), 10.4% (.35), and 0.79 (.4) for, respectively, canopy height, cover, and plant area index, and, matching the mapping quality of supervised methods (Figures I1 and I2). For tree species identification ($PF$), DUNIA slightly underperforms by 0.8% at KNN=50 but exceeds the baseline by 1.4% when decreasing the number of neighbors (KNN=5) used for the distance-weighted voting. For land cover mapping ($CLC_{+}$), DUNIA achieves strong performance with a wF1 score of 80.1%. However, for crop type mapping ($PASTIS$), DUNIA significantly underperforms compared to the state-of-the-art, which is likely due to the variability of the phenological cycles of crops that cannot be well captured by a single median composite.
@@
For lower KNN values, the performance of vertical structure-related products and tree species identification remains stable. In contrast, land cover and crop type mapping experience significant performance drops, indicating that the model needs more samples for accurate classification. For retrieval databases with fewer samples (Table G1), vertical structure-related products ($\mathcal{W}_{rh}$, $\mathcal{W}_{c}$, $\mathcal{W}_{pai}$) maintain their performance and map qualities (Figure I4), while other products ($CLC_{+}$, $PASTIS$, $PF$) show a substantial degradation in performance.
@@
Regarding waveform ($\mathcal{W}$) retrieval performance, we observe that our model is able to retrieve relevant waveforms given a pixel embedding as a query, with a correlation coefficient of .70 for S=50k l, and decreases by a small factor with smaller retrieval database sizes (Table G1). This is to be expected given the small subset of waveforms in the retrieval database that might not reflect all possible variations for a particular height class. Retrieved vs. reference waveforms can be found in Figures I5, I6, I7 and I8.
4.2.2 Fine-tuning Performance
Table 2: Fine-tuning performance of DUNIA and the five competing models. and are DUNIA’s embeddings from $\mathcal{O^{V}}$ and $\mathcal{O^{H}}$ respectively. S represents the number of samples used for the fine-tuning, with im meaning a $64\times 64$ pixels fully annotated image, and l meaning a single annotated pixel. $\mathcal{W}^{**}$ represents performance results for vertical structures higher than 5 m. Best scores are in bold.
Dataset Metric Samples (S) DUNIA AnySat Croma DOFA DeCUR SatMAE
$\mathcal{W}_{rh}$ RMSE (r) 5K im 1.3 (.95) 2.8 (.89) 3.5 (.78) 11.0 (.51) 11.0 (.55) 10.5 (.52)
$\mathcal{W}_{c}$ RMSE (r) 5K im 9.8 (.85) 12.1 (.79) 14.2 (.73) 29.2 (.50) 28.5 (.48) 30.2 (.48)
$\mathcal{W}_{pai}$ RMSE (r) 5K im 0.62 (.71) 0.95 (.67) 1.2 (.61) 1.5 (.38) 1.6 (.38) 1.6 (.39)
$CLC_{+}$ wF1 5K im 90.3 90.1 86.4 72.0 75.1 75.0
$PASTIS$ wF1 1.5K im 77.0 81.1 73.3 54.5 57.3 55.2
$PF$ wF1 50K l 82.2 82.3 80.5 79.8 78.9 78.8
$\mathcal{W}^{**}$ r $\approx$19M l .78